• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 04:52:50 on January21,2008

    Entry number 226526

    Followups:

    keyword=PREX stress test. part 1.

    Thick PREX target.
    (X,Y) = (0.1) = nominally good alignment.

    Raster = 6.5 x 5 (MCC setting) = 4 x 4 (reality) = 2.7 x 2.7 (obs.)
    (see previous halogs for explanation)

    Will ramp the beam current and monitor rates and raster XY.

    Run 3089 start at 3:51
    This run will go continuously.

    Time ... Current ... T1 .... T3 .... Comments

    03:51 ... 2 uA ..... 1030 ... 640

    03:56 ... 10 uA .... 5300 ... 3300

    04:08 ... MCC stops for a QE measurement.
    Also need to reset a "box" of some sort ...

    04:20 beam returns. We have asked for 20 uA

    04:20 .... 20 uA ... 10.75 kHz ... 6400 Hz

    04:30 .... 30 uA ... 15.3 k ...... 9.45 K

    Wanted to go to 40 uA. Cannot because of ion chambers tripping.
    MCC will call Radcon to get permission.

    Remain at 30 uA for now.




    FIGURE 1